# CurvFed: Curvature-Aligned Federated Learning for Fairness without Demographics

### v6

@@ 5.2. Evaluation on CurvFed’s Effectiveness
To establish the effectiveness of CurvFed, we have investigated the following research questions (RQs), each targeting a different aspect of fairness and efficacy in FL. The results addressing these questions are presented in corresponding evaluation tables as indicated below.
@@
We compare CurvFed against standard benign FL models—FedAvg, FedSAM, and FedSWA—and two recent fairness-without-disparity (FWD) approaches: KD-FedAvg (chai2022fairness, ) (adopted for FL settings) and FedSRCVaR (papadaki2024federated, ). Fairness is evaluated using the FATE score and Equal Opportunity (EO) gap with respect to sex.
Dataset Model F1 Score EO Gap FATE Score $\times 10^{-1}$
WIDAR FedAvg 0.85 0.4263 0.0000
FedSAM 0.8005 0.4251 -0.56
FedSWA 0.8405 0.4071 0.33
FedSRCVaR 0.7775 0.3828 0.160
KD-Fedavg 0.8279 0.4097 0.122
CurvFed 0.8100 0.3533 1.233
Stress sensing FedAvg 0.7253 0.3069 0.0000
FedSAM 0.7758 0.3034 0.812
FedSWA 0.7406 0.3015 0.384
FedSRCVaR 0.6692 0.2734 0.319
KD-Fedavg 0.7154 0.2522 1.645
CurvFed 0.7814 0.2522 2.559
HugaDB FedAvg 0.8496 0.4369 0.0000
FedSAM 0.8555 0.4255 0.329
FedSWA 0.8418 0.4242 0.1917
FedSRCVaR 0.8354 0.3939 0.815
KD-Fedavg 0.8342 0.4166 0.281
CurvFed 0.8855 0.3952 1.375
Table 1. Fairness–efficacy balance comparisons for RQ-1
FedAvg is the fairness-unaware baseline, while FedSAM and FedSWA are benign baselines that promote flatter loss landscapes via client-side SAM and server-side SWA (caldarola2022improving, ).
@@
Evaluation discussion: As shown in Table 1, CurvFed consistently improves the fairness efficacy tradeoff. In WIDAR, it reduces the EO gap from 0.4263 (FedAvg) to 0.3533 with minimal F1 reduction (0.8100 vs. 0.8500). In Stress Sensing, it improves both F1 (0.7814 vs. 0.7253) and EO gap (0.2522 vs. 0.3069). HugaDB shows similar trends, with a higher F1 (0.8855) and reduced EO gap (0.3952) compared to FedAvg.
CurvFed consistently outperforms KD-FedAvg and FedSRCVaR in FATE scores across all datasets—Stress Sensing (2.559 vs. 1.645 and 0.319), WIDAR (1.233 vs. 0.122 and 0.160), and HugaDB (1.375 vs. 0.281 and 0.815). While KD-based methods excel in centralized settings with diverse group-wise data (chai2022fairness, ), their local fairness improvements do not reliably carry over after model aggregation in FL (hamman2023demystifying, ). This limits their effectiveness in FL settings.
FedSRCVaR, relies on a constraint $\rho$ to bound worst-case group sizes (papadaki2024federated, ), assumes known group distributions—often unavailable in human-centric data with missing or highly variable groups across clients. This leads to unstable optimization and poor fairness–utility trade-off (e.g., F1 drops to 0.6692 in the Stress Sensing dataset). In contrast, CurvFed makes no such assumptions and achieves better fairness–performance balance.
5.2.2. RQ-2: Can fairness be achieved in FL settings that include both single-person and multi-person clients?.
Traditional fairness-aware FL methods assume each client contains data from all sensitive groups. However, as noted in Section 1 challenge 3, edge devices may contain data from a single user or multiple users. Thus, this assumption often fails in real-world deployments. Notably, CurvFed removes this assumption by promoting fairness through loss landscape regularization, making it impactful for practical deployment in human-sensing applications. Table 2 presents CurvFed’s performance across all three datasets, under a mix of single- and multi-person clients with 1–2 males and 0–1 females per client (See Appendix for detailed client distributions). Despite the absence of full group representation, CurvFed consistently outperforms other FL methods, achieving strong FATE scores of $0.770\times 10^{-1}$ (WIDAR), $4.604\times 10^{-1}$ (Stress Sensing), and $0.314\times 10^{-1}$ (HugaDB), demonstrating balanced fairness and performance under practical constraints.
@@ 5.2.3. RQ-3: Can CurvFed enhance fairness specifically for single-person clients?
Only single-person clients are common in human-centric FL but pose challenges due to the lack of intra-client diversity, making global fairness and performance difficult to achieve. Prior studies (lu2023fedcom, ; yao2024rethinking, ) show that simple aggregation methods like FedAvg often degrade global performance in these cases (lu2023fedcom, ).
We evaluated CurvFed against other FWD baselines in only single-client scenarios. As shown in Table 2, CurvFed consistently achieves the best fairness–accuracy trade-off, with the highest FATE scores across all datasets: $0.601\times 10^{-1}$ for WIDAR, $3.511\times 10^{-1}$ for Stress Sensing, and $0.348\times 10^{-1}$ for HugaDB —demonstrating its effectiveness even without client-level sensitive attribute diversity.
@@
Dataset Client Type Model F1 Score EO Gap FATE Score $\times 10^{-1}$
WIDAR Single & Multi-Person FedAvg 0.8361 0.4289 0.0000
FedSAM 0.8311 0.4122 0.329
FedSWA 0.8540 0.4366 0.0364
KD-FedAvg 0.7844 0.4085 -0.140
FedSRCVaR 0.7105 0.4085 -1.024
CurvFed 0.7938 0.3742 0.770
Single Person Only FedAvg 0.8420 0.4302 0.0000
FedSAM 0.8383 0.4148 0.313
FedSWA 0.8421 0.4289 0.0321
FedSRCVaR 0.7064 0.4149 -1.253
KD-FedAvg 0.7691 0.4097 -0.389
CurvFed 0.7998 0.3828 0.601
Stress Sensing Single & Multi-Person FedAvg 0.7542 0.3774 0.0000
FedSAM 0.7467 0.3877 -0.371
FedSWA 0.8350 0.3121 2.802
KD-FedAvg 0.7471 0.2487 3.318
FedSRCVaR 0.6967 0.2751 1.948
CurvFed 0.8300 0.2416 4.604
Single Person Only FedAvg 0.7607 0.3827 0.0000
FedSAM 0.7663 0.3395 1.204
FedSWA 0.7820 0.3633 0.786
FedSRCVaR 0.6757 0.2222 3.076
KD-FedAvg 0.6764 0.2522 2.303
CurvFed 0.7298 0.2328 3.511
HugaDB Single & Multi-Person FedAvg 0.8610 0.4380 0.0000
FedSAM 0.8480 0.4142 0.293
FedSWA 0.8404 0.4107 0.287
KD-FedAvg 0.7815 0.4119 -0.537
FedSRCVaR 0.8271 0.4059 0.236
CurvFed 0.8163 0.3976 0.314
Single Person Only FedAvg 0.8720 0.4542 0.0000
FedSAM 0.8565 0.4980 -1.142
FedSWA 0.7835 0.4250 -0.372
FedSRCVaR 0.7898 0.4123 -0.020
KD-FedAvg 0.7797 0.4174 -0.248
CurvFed 0.8317 0.4174 0.348
Table 2. Performance comparison across different client setups: single & multi-person clients (RQ2) and single-person clients (RQ3)
@@ 5.2.4. RQ-4: Could CurvFed improve fairness across multiple sensitive attributes simultaneously?
Dataset Sensitive Model F1 Score EO Gap FATE Score $\times 10^{-1}$
Attribute
WIDAR FedAvg 0.8436 0.5130 0.0000
Orientation FedSAM 0.8007 0.5082 -0.414
FedSWA 0.8494 0.5146 0.037
FedSRCVaR 0.7772 0.4318 0.795
KD-Fedavg 0.8279 0.4756 0.543
CurvFed 0.8100 0.3814 2.154
Stress Sensing FedAvg 0.7253 0.4797 0.0000
Hand FedSAM 0.7758 0.4268 1.799
FedSWA 0.7406 0.4638 0.928
FedSRCVaR 0.6692 0.3651 1.617
KD-Fedavg 0.7154 0.3957 1.615
CurvFed 0.7814 0.4320 1.766
Table 3. Performance metrics comparison under different sensitive attributes for RQ4
@@
In FL-based human sensing, data often simultaneously involve multiple sensitive attributes (e.g., sex, age, body type, sensor placement). To assess CurvFed under such conditions, we evaluated fairness with respect to additional sensitive attributes—hands’ (watch-hand) and orientation’ (sensor placement)—in the Stress Sensing and WIDAR datasets, using the same models as in Table 1. For WIDAR, we induced disparity by subsampling 50% of data from orientations 4 and 5, grouping orientations 1–3 as "major" and the rest as "minor." In Stress Sensing, only 13 of 48 participants had both-hand data, while others had left-hand only. HugaDB, with only sex as a sensitive attribute, was excluded.
As shown in Table 3, CurvFed achieves strong FATE scores across sensitive attributes, underscoring its effectiveness in enhancing fairness while maintaining competitive accuracy. Although FedSAM slightly outperforms others in Stress Sensing with CurvFed being second best. Notably, both the approaches target the loss landscape curvature attributes (foret2021sharpnessaware, ). CurvFed delivers the highest FATE score on WIDAR. These results highlight CurvFed’s capability to effectively balance fairness and accuracy, thereby addressing Challenge 4 from Section 1.
5.3. Feasibility Study
This section presents a comprehensive set of experiments to evaluate the real-world feasibility of CurvFed.
5.3.1. Feasibility of Real-World Deployment with CurvFed:
Client Device
Training
Time (s)
Memory
Train
Usage (%)
Upload
Time (s)
Client 1 Google Pixel 6 0.7819 268.48 27.65 3.48
Client 2 Google Pixel 6 0.525 267.88 28.63 4.03
Client 3 Raspberry Pi 5 0.4732 281.36 41.92 3.20
Client 4 Intel i5-14600T 0.206 256.80 3.34 3.39
Client 5 Intel i5-14600T 0.239 259.69 2.45 3.48
Client 6 Jetson Nano 0.409 288.27 7.54 1.903
Table 4. Per-client training statistics from the real-world FL deployment. Metrics include training time, Process RSS memory usage (MB), average CPU utilization, and model upload latency.
To evaluate the real-world deployability of CurvFed, we conducted a live FL deployment using a heterogeneous set of six edge devices: two Google Pixel 6 smartphones, a Raspberry Pi 5, an NVIDIA Jetson Nano, and two Intel i5-14600T desktops (Fig. 2). Each device was assigned data from a unique participant in the Stress Sensing dataset (Section 5.1) and exchanged model updates with a central server via Google Cloud APIs authenticated through PyDrive.

### v7

@@ 5.3. Evaluation on CurvFed’s Effectiveness
To establish the effectiveness and stability of CurvFed, we conducted a comprehensive evaluation using stratified 5-fold cross-validation (maintaining an 80:20 train-test split per client), repeated across three distinct random seeds. To maintain clarity, we present the mean and standard deviation for both the F1 score and the EO gap, averaged across folds for a representative seed. Furthermore, we report the FATE score, which is calculated using the mean F1 and mean EO values as defined in Equation 2. Complete results for the additional seeds are provided in Appendix A.1.4. Our analysis is structured around the following research questions (RQs), each targeting a different aspect of fairness and efficacy in FL:
@@
We evaluated whether CurvFed improves the trade-off between predictive performance and fairness relative to existing federated learning baselines. Fairness is measured using the FATE score and the Equal Opportunity (EO) gap with respect to sex.
Dataset Model F1 Score Mean $\pm$ Std ($\uparrow$) EO Gap Mean $\pm$ Std ($\downarrow$) FATE $\times 10^{-1}$ Mean Score ($\uparrow$)
WIDAR FedAvg $\textbf{0.862}\pm 0.0006$ $412\pm 0.0154 0.0$0000.0000
FedSAM $0.843\pm 0.0108$ $401\pm 0.0230 0.0$560.056
FedSWA $0.859\pm 0.0030$ $0.405\pm 0.0142$ $0.147$
FedSRCVaR $0.724\pm 0.0187$ $345\pm 0.0534 0.0$190.019
KD-Fedavg $0.849\pm 0.0123$ $0.392\pm 0.0299$ $0.327$
CurvFed $0.806\pm 0.0458$ $\textbf{0.354}\pm 0.0099$ 0.741
Stress sensing FedAvg $0.767\pm 0.0146$ $3240\pm 0.014 0.0$$0$
FedSAM $0.784\pm 0.0094$ $0.3135\pm 0.0103$ $0.548$
FedSWA $0.743\pm 0.0164$ $0.305\pm 0.0034$ $0.257$
FedSRCVaR $0.694\pm 0.0253$ $0.278\pm 0.0231$ $0.458$
KD-Fedavg $0.717\pm 0.0157$ $0.259\pm 0.0095$ $1.344$
CurvFed $\textbf{0.789}\pm 0.0086$ $\textbf{0.259}\pm 0.0045$ 2.292
HugaDB FedAvg $0.852\pm 0.0063$ $482\pm 0.0294 0.0$000
FedSAM $0.786\pm 0.0034$ $438\pm 0.0042 0.0$140.014
FedSWA $0.809\pm 0.0033$ $439\pm 0.0144 0.0$400.040
FedSRCVaR $0.862\pm 0.0032$ $0.410\pm 0.0103$ $0.163$
KD-Fedavg $0.776\pm 0.0061$ $444\pm 0.0161 0.0$08-0.008
CurvFed $\mathbf{0.892}\pm 0.0072$ $\mathbf{0.406}\pm 0.0042$ $\mathbf{0.206}$
Percept-R FedAvg 0.7774 $\pm$ 0.002 0.162 $\pm$ 0.005 0.0000
FedSAM 0.777 $\pm$0.002 0.136 $\pm$ 0.01 1.72
FedSWA 0.777 $\pm$ 0.002 0.127 $\pm$ 0.01 1.66
FedSRCVaR 0.771 $\pm$ 0.004 0.130 $\pm$ 0.01 1.73
KD-Fedavg 0.777 $\pm$ 0.001 0.150 $\pm$ 0.01 0.76
CurvFed 0.764 $\pm$ 0.003 0.106 $\pm$ 0.01 3.29
Table 3. Fairness–efficacy balance comparisons for RQ-1. Results are averaged over 5-fold cross-validation. Metrics with $\uparrow$ indicate higher is better, while metrics with $\downarrow$ indicate lower is better. CurvFed attains the highest FATE score, reflecting the overall fairness–utility balance
FedAvg is the fairness-unaware baseline, while FedSAM and FedSWA are benign baselines that promote flatter loss landscapes via client-side SAM and server-side SWA (Caldarola et al., 2022).
@@
Evaluation discussion: As shown in Table 3, results are reported as the mean $\pm$ standard deviation over 5-fold cross-validation. CurvFed consistently improves the fairness–efficacy trade-off across all datasets. In WIDAR, it reduces the EO gap from $0.412\pm 0.0154$ (FedAvg) to $0.354\pm 0.0099$, with a moderate F1 reduction ($0.806\pm 0.0458$ vs. $0.862\pm 0.0006$). In Stress Sensing, it improves both F1 ($0.789\pm 0.0086$ vs. $0.767\pm 0.0146$) and EO gap ($0.259\pm 0.0045$ vs. $0.324\pm 0.014$). HugaDB shows similar trends, with a higher F1 ($0.892\pm 0.0072$) and reduced EO gap ($0.406\pm 0.0042$) compared to FedAvg ($0.852\pm 0.0063$, $0.482\pm 0.0294$). In Percept-R, where baseline models exhibit comparable F1 performance, CurvFed achieves the lowest EO gap ($0.106\pm 0.01$) with only a small reduction in F1 ($0.764\pm 0.003$ vs. $0.777\pm 0.002$ for FedAvg).
CurvFed consistently outperforms KD-FedAvg and FedSRCVaR in FATE scores across all datasets—Stress Sensing ($2.292$ vs. $1.344$ and $0.458$), WIDAR ($0.741$ vs. $0.327$ and $0.019$), HugaDB ($0.206$ vs. 0.008-0.008 and $0.163$), and Percept-R ($3.29$ vs. $0.76$ and $1.73$). While KD-based methods excel in centralized settings with diverse group-wise data (Chai et al., 2022), their local fairness improvements do not reliably carry over after model aggregation in FL (Hamman and Dutta, 2023). This limits their effectiveness in FL settings.
FedSRCVaR, relies on a constraint $\rho$ to bound worst-case group sizes (Papadaki et al., 2024), assumes known group distributions—often unavailable in human-centric data with missing or highly variable groups across clients. This results in unstable optimization and poor fairness–utility trade-offs, as reflected by the reduced F1 score ($0.694\pm 0.0253$) on the Stress Sensing dataset. In contrast, CurvFed makes no such assumptions and achieves better fairness–performance balance.
5.3.2. RQ-2: Can fairness be achieved in FL settings that include both single-person and multi-person clients?
Traditional fairness-aware FL methods assume each client contains data from all sensitive groups. However, as noted in Section 1 challenge 3, edge devices may contain data from a single user or multiple users. Thus, this assumption often fails in real-world deployments. Notably, CurvFed removes this assumption by promoting fairness through loss landscape regularization, making it impactful for practical deployment in human-sensing applications. Table 4 presents CurvFed’s performance across all four datasets, under a mix of single- and multi-person clients with 1–2 males and 0–1 females per client (See Appendix A.1.3 for detailed client distributions). Despite the absence of full group representation, CurvFed consistently outperforms other FL methods, achieving strong FATE scores of $0.479\times 10^{-1}$ on WIDAR, $4.032\times 10^{-1}$ on Stress Sensing, $0.590\times 10^{-1}$ on HugaDB, and $2.03\times 10^{-1}$ on Percept-R, demonstrating balanced fairness and performance under practical constraints.
@@ 5.3.3. RQ-3: Can CurvFed enhance fairness specifically for single-person clients?
Only single-person clients are common in human-centric FL but pose challenges due to the lack of intra-client diversity, making global fairness and performance difficult to achieve. Prior studies (Lu et al., 2023; Yao et al., 2024) show that simple aggregation methods like FedAvg often degrade global performance in these cases (Lu et al., 2023).
@@
Dataset Client Type Model F1 Score Mean $\pm$ Std ($\uparrow$) EO Gap Mean $\pm$ Std ($\downarrow$) FATE $\times 10^{-1}$ Mean ($\uparrow$)
WIDAR Single & Multi-Person FedAvg $0.836\pm 0.0132$ $429\pm 0.0081 0.0$000.000
FedSAM $0.822\pm 0.0143$ $0.411\pm 0.0032$ $0.238$
FedSWA $\mathbf{0.852}\pm 0.0323$ $0.431\pm 0.0034$ $0.140$
KD-FedAvg $0.781\pm 0.0030$ $0.389\pm 0.0013$ $0.270$
FedSRCVaR $0.712\pm 0.0048$ $0.405\pm 0.0112$ 0.929-0.929
CurvFed $0.791\pm 0.0089$ $\mathbf{0.385}\pm 0.0089$ $\mathbf{0.479}$
Single Person Only FedAvg $\mathbf{0.843}\pm 0.0115$ $441\pm 0.0143 0.0$000.000
FedSAM $0.839\pm 0.0053$ $0.423\pm 0.0098$ $0.366$
FedSWA $0.842\pm 0.0093$ $0.424\pm 0.0034$ $0.383$
FedSRCVaR $0.706\pm 0.0124$ $0.418\pm 0.0075$ 1.094-1.094
KD-FedAvg $0.771\pm 0.0119$ $402\pm 0.0082 0.0$300.030
CurvFed $0.793\pm 0.0092$ $\mathbf{0.382}\pm 0.0045$ $\mathbf{0.754}$
Stress Sensing Single & Multi-Person FedAvg $0.761\pm 0.0143$ $376\pm 0.0023 0.0$000.000
FedSAM $0.749\pm 0.0034$ $0.382\pm 0.0120$ 0.313-0.313
FedSWA $\mathbf{0.841}\pm 0.0145$ $0.310\pm 0.0114$ $2.822$
KD-FedAvg $0.745\pm 0.0129$ $0.248\pm 0.0024$ $3.193$
FedSRCVaR $0.701\pm 0.0083$ $0.283\pm 0.0215$ $1.683$
CurvFed $0.829\pm 0.0129$ $\mathbf{0.258}\pm 0.0134$ $\mathbf{4.032}$
Single Person Only FedAvg $0.762\pm 0.0014$ $382\pm 0.0034 0.0$000.000
FedSAM $0.769\pm 0.0093$ $0.338\pm 0.0093$ $1.235$
FedSWA $0.779\pm 0.0182$ $0.368\pm 0.0076$ $0.589$
FedSRCVaR $0.675\pm 0.0214$ $0.218\pm 0.0142$ $3.143$
KD-FedAvg $0.680\pm 0.0313$ $0.257\pm 0.0117$ $2.186$
CurvFed $\mathbf{0.731}\pm 0.0362$ $\mathbf{0.235}\pm 0.0189$ $\mathbf{3.447}$
HugaDB Single & Multi-Person FedAvg $\mathbf{0.856}\pm 0.0113$ $441\pm 0.0042 0.0$000.000
FedSAM $0.847\pm 0.0118$ $0.408\pm 0.0032$ $0.636$
FedSWA $0.842\pm 0.0092$ $0.413\pm 0.0092$ $0.461$
KD-FedAvg $0.789\pm 0.0068$ $409\pm 0.0045 0.0$48-0.048
FedSRCVaR $0.827\pm 0.0182$ $0.403\pm 0.0142$ $0.524$
CurvFed $0.816\pm 0.0340$ $\mathbf{0.395}\pm 0.0083$ $\mathbf{0.590}$
Single Person Only FedAvg $\mathbf{0.871}\pm 0.0013$ $454\pm 0.0034 0.0$000.000
FedSAM $0.856\pm 0.0052$ $0.502\pm 0.0182$ 1.230-1.230
FedSWA $0.783\pm 0.0432$ $0.432\pm 0.0231$ 0.525-0.525
FedSRCVaR $0.792\pm 0.0092$ $411\pm 0.0092 0.0$380.038
KD-FedAvg $0.769\pm 0.0082$ $\mathbf{0.409}\pm 0.0048$ 0.180-0.180
CurvFed $0.824\pm 0.0054$ $0.418\pm 0.0092$ $\mathbf{0.253}$
Percept-R Single & Multi-Person FedAvg 0.785 $\pm$ 0.006 0.132 $\pm$ 0.01 0.0000
FedSAM 0.781 $\pm$ 0.004 0.113 $\pm$ 0.01 1.31
FedSWA 0.780 $\pm$ 0.004 0.117 $\pm$ 0.01 1.02
KD-FedAvg 0.784 $\pm$ 0.006 0.132 $\pm$ 0.01 -0.06
FedSRCVaR 0.783 $\pm$ 0.006 0.131 $\pm$ 0.009 0.03
CurvFed 0.772 $\pm$ 0.003 0.102 $\pm$ 0.01 2.03
Single Person Only FedAvg 0.770 $\pm$ 0.005 0.108 $\pm$ 0.01 0.0000
FedSAM 0.768 $\pm$ 0.006 0.115 $\pm$ 0.008 -0.71
FedSWA 0.765 $\pm$ 0.007 0.112 $\pm$ 0.008 -0.45
FedSRCVaR 0.770 $\pm$ 0.005 0.102 $\pm$ 0.01 0.5
KD-FedAvg 0.770 $\pm$ 0.005 0.106 $\pm$ 0.01 0.17
CurvFed 0.764 $\pm$ 0.007 0.054 $\pm$ 0.004 4.85
Table 4. Performance comparison across different client setups: single & multi-person clients (RQ2) and single-person clients (RQ3). Results are averaged over 5-fold cross-validation. Metrics with $\uparrow$ indicate higher is better, while metrics with $\downarrow$ indicate lower is better. Note: While some methods perform better on individual F1 or EO Gap metrics, CurvFed attains the highest FATE score, reflecting the overall fairness–utility balance
We evaluated CurvFed against other FWD baselines in only single-client scenarios. As shown in Table 4, CurvFed consistently achieves the highest FATE scores across all datasets in this setting, including $0.754\times 10^{-1}$ for WIDAR, $3.447\times 10^{-1}$ for Stress Sensing, $0.253\times 10^{-1}$ for HugaDB, and $4.85\times 10^{-1}$ for Percept-R —demonstrating its effectiveness even without client-level sensitive attribute diversity.
@@ 5.3.4. RQ-4: Could CurvFed improve fairness across multiple sensitive attributes simultaneously?
In FL-based human sensing, data often simultaneously involve multiple sensitive attributes (e.g., sex, age, body type, sensor placement). To assess CurvFed under such conditions, we evaluated fairness with respect to additional sensitive attributes—hands’ (watch-hand) and orientation’ (sensor placement)—in the Stress Sensing and WIDAR datasets, using the same models as in Table 3. For WIDAR, we induced disparity by subsampling 50% of data from orientations 4 and 5, grouping orientations 1–3 as "major" and the rest as "minor." In Stress Sensing, only 13 of 48 participants had both-hand data, while others had left-hand only. HugaDB and Percept_R, with only sex as a sensitive attribute, were excluded.
As shown in Table 5, CurvFed consistently achieves the strongest fairness–utility trade-off across both sensitive attributes. Although other methods may attain higher F1 scores or lower EO gaps individually, CurvFed maintains competitive classification performance while simultaneously reducing EO gaps, resulting in the highest FATE scores on both WIDAR and Stress Sensing. This indicates that CurvFed does not optimize a single objective in isolation but instead preserves a balanced trade-off between fairness and accuracy when multiple sensitive attributes are present, thereby addressing Challenge 4 outlined in Section 1.
Dataset Sensitive Attribute Model F1 Score Mean $\pm$ Std ($\uparrow$) EO Gap Mean $\pm$ Std ($\downarrow$) FATE $\times 10^{-1}$ Mean ($\uparrow$)
WIDAR Orientation FedAvg $\mathbf{0.862}\pm 0.0006$ $506\pm 0.0161 0.0$000.000
FedSAM $0.843\pm 0.0108$ $0.481\pm 0.0146$ $0.284$
FedSWA $0.859\pm 0.0030$ $0.499\pm 0.0499$ $0.102$
FedSRCVaR $0.725\pm 0.0187$ $422\pm 0.0535 0.0$600.060
KD-Fedavg $0.849\pm 0.0123$ $0.458\pm 0.0343$ $0.793$
CurvFed $0.806\pm 0.0458$ $\mathbf{0.395}\pm 0.0507$ $\mathbf{1.550}$
Stress Sensing Hand FedAvg $0.767\pm 0.0146$ $464\pm 0.0124 0.0$000.000
FedSAM $0.785\pm 0.0094$ $0.463\pm 0.0830$ $0.233$
FedSWA $0.744\pm 0.0164$ $0.441\pm 0.0234$ $0.173$
FedSRCVaR $0.695\pm 0.0253$ $\mathbf{0.394}\pm 0.0211$ $0.550$
KD-Fedavg $0.718\pm 0.0157$ $0.406\pm 0.0193$ $0.587$
CurvFed $\mathbf{0.789}\pm 0.0086$ $0.433\pm 0.0023$ $\mathbf{0.945}$
Table 5. Fairness–efficacy balance under different sensitive attributes for RQ-4.Results are averaged over 5-fold cross-validation. Metrics with $\uparrow$ indicate higher is better, while metrics with $\downarrow$ indicate lower is better. Note: While some methods perform better on individual F1 or EO Gap metrics, CurvFed attains the highest FATE score, reflecting the overall fairness–utility balance
6. Feasibility Study
In this section, we conduct a rigorous empirical study of CurvFed in real-world settings. We systematically evaluate its deployment feasibility (Section 6.1) with additional runtime benchmarking/overhead analysis detailed in Appendix B. Furthermore, we discuss sensitivity to key factors (Sections 6.2 and 6.2.4), and conclude with an ablation study that substantiates the effectiveness of the core design choices (Section 6.3).
6.1. Assessing Real-World Deployment Feasibility of CurvFed’s Modules